# ECORE: Energy-Conscious Optimized Routing for Deep Learning Models at the Edge

### v3

@@ V Related Work
TABLE III: Comparison of Related Work and ECORE
Study Edge Com. Obj. Det. Dynamic Routing Energy Cons. Accuracy Real Testbed
Ji et al. [8]
Trinh et al. [9]
Tu et al. [10]
Zhang et al. [11]
Tundo et al. [12]
Matathammal et al. [13]
Kulkarni et al. [14]
Marda et al. [15]
Stripelis et al. [16]
Maurya et al. [17]
Zheng et al. [18]
Guha et al. [19]
Mohammadshahi et al. [20]
Sikeridis et al. [21]
Jiang et al. [22]
Kang et al. [23]
Davis and Belviranli [24]
Zhang et al. [25]
ECORE
This section surveys related studies and situates our approach (see Table III). We organize prior work into three scopes: (i) dynamic on-device adaptation, (ii) system-level orchestration across edge clusters and between edge and cloud, and (iii) routing policies. The first includes pipeline reconfiguration, local model and accelerator selection, and contextual-bandit policies within a single device. The second covers multi-stream configuration, bandwidth allocation, cluster scheduling, and offloading across edge and cloud. The third reviews query-aware dispatch among model pools in the cloud, which is conceptually related but outside our edge-vision scope.
@@ V-B System-level orchestration
A second line of research focuses on coordinating resources across different nodes and tiers. EdgeLeague [10] and JCAB [11] adapt detector configurations and uplink bandwidth for multi-stream analytics under fluctuating network conditions, scheduling resources to meet quality-of-service and latency requirements. VideoStorm [25] performs cluster-scale scheduling based on resource and quality profiles, emphasizing throughput and deadline compliance rather than per-image content decisions. In cloud–edge offloading, CELTC [8] coordinates large cloud models with lightweight edge models using deep reinforcement learning, while Trinh et al. [9] determine when to offload vision tasks under energy constraints. These orchestration approaches optimize configurations, bandwidth, and scheduling across the system, whereas our work operates entirely at the edge and makes content-aware selections of the most suitable device–model pair for each image.
@@ VI Conclusions and Future Direction
Deploying object detection models on resource-constrained edge devices, especially those powered by batteries or renewable energy sources such as solar panels, demands careful consideration of energy consumption. In many real-world applications, particularly safety-critical systems, the environment is highly dynamic and the number of objects captured in each frame can vary greatly. This variability directly impacts the accuracy of object detection models and highlights the need for adaptive strategies that maintain a balance between efficiency and performance. To address these challenges, we introduce ECORE together with three dynamic, estimation-based routing algorithms. Each algorithm begins by applying lightweight preprocessing to estimate the number of objects in an incoming image request. Based on this estimate, the request is then routed to the most suitable edge model–device pair, a decision informed by detailed profiling of accuracy and energy trade-offs. This design aims to minimize overall energy consumption while ensuring that detection accuracy remains within an acceptable range. Through extensive experimental evaluation, we show that the proposed strategy consistently outperforms baseline approaches. In particular, it achieves significant reductions in energy usage without compromising accuracy, thereby demonstrating the effectiveness of dynamic, context-aware routing as a promising direction for enabling energy-efficient AI workloads at the edge.

### v4

@@ V Related Work
This section surveys related studies and situates our approach. We organize prior work into three scopes: (i) dynamic on-device adaptation, (ii) system-level orchestration across edge clusters and between edge and cloud, and (iii) routing policies. The first includes pipeline reconfiguration, local model and accelerator selection, and contextual-bandit policies within a single device. The second covers multi-stream configuration, bandwidth allocation, cluster scheduling, and offloading across edge and cloud. The third reviews query-aware dispatch among model pools in the cloud, which is conceptually related but outside our edge-vision scope.
@@ V-B System-level orchestration
A second line of research focuses on coordinating resources across different nodes and tiers. EdgeLeague [22] and JCAB [25] adapt detector configurations and uplink bandwidth for multi-stream analytics under fluctuating network conditions, scheduling resources to meet QoS and latency requirements. VideoStorm [24] performs cluster-scale scheduling based on resource and quality profiles, emphasizing throughput and deadline compliance rather than per-image content decisions. In cloud–edge offloading, CELTC [6] coordinates large cloud models with lightweight edge models using deep reinforcement learning, while Trinh et al. [21] determine when to offload vision tasks under energy constraints. These orchestration approaches optimize configurations, bandwidth, and scheduling across the system, whereas our work operates entirely at the edge and makes content-aware selections of the most suitable device–model pair for each image.
@@ VI Conclusions and Future Direction
Deploying object detection models on resource-constrained edge devices, especially those powered by batteries or renewable energy sources such as solar panels, demands careful consideration of energy consumption. In many real-world applications, particularly safety-critical systems, the environment is highly dynamic and the number of objects captured in each frame can vary greatly. This variability directly impacts the accuracy of object detection models and highlights the need for adaptive strategies that maintain a balance between efficiency and performance. To address these challenges, we introduce ECORE together with three dynamic, estimation-based routing algorithms. Each algorithm begins by applying lightweight preprocessing to estimate the number of objects in an incoming image request. Based on this estimate, the request is then routed to the most suitable edge device-model pair, a decision informed by detailed profiling of accuracy and energy trade-offs. This design aims to minimize overall energy consumption while ensuring that detection accuracy remains within an acceptable range. Through extensive experimental evaluation, we show that the proposed strategy consistently outperforms baseline approaches. In particular, it achieves significant reductions in energy usage without compromising accuracy, thereby demonstrating the effectiveness of dynamic, context-aware routing as a promising direction for enabling energy-efficient AI workloads at the edge.
@@
Acknowledgment
We would like to acknowledge support through the Australian Research Council’s funded projects DP230100081 and LP210200213.